OpenStudy (anonymous):

Find the amplitude and period of f (x) = −2 cos(1/3x + 1

OpenStudy (mertsj):

the amplitude is the coefficient of the function, that is 2

OpenStudy (mertsj):

rewrite cos(1/3x + 1)=cos1/3[(x) +3]

OpenStudy (mertsj):

Then the period will be 2pi/1/3 or 6 pi

OpenStudy (anonymous):

thanks can you help Find the angle co-terminal with 23pi/3

OpenStudy (mertsj):

23pi/3 = 7 and 2/3 pi. Since each revolution is 2 pi, we make 3 complete revolutions and then 1 2/3 pi more or 5/3 pi. So the angle that is co-terminal with 23pi/3 is 5/3 pi

OpenStudy (anonymous):

i dont have 5pi/3 as an option

OpenStudy (mertsj):

It could be - pi/3

OpenStudy (anonymous):

yeah i have that option

OpenStudy (mertsj):

-pi/3 is the same as 5pi/3

OpenStudy (anonymous):

how do you convert from 5pi/3 to -1pi/3

OpenStudy (mertsj):

5pi/3 - 2pi

OpenStudy (anonymous):

ok i see

OpenStudy (mertsj):

Good

OpenStudy (anonymous):

you think you can help me with this one too? Find the exact value of cot (sin^-1 (sqrt2/2))

OpenStudy (mertsj):

(sin^-1 (sqrt2/2) meant the angle whose sin is sqrt2/2 which is pi/4 or 3pi/4

OpenStudy (mertsj):

so now the problem becomes cot(pi/4) which is 1 or cot(3pi/4) which is -1.

OpenStudy (mertsj):

You might want to reread your directions carefully. If you have to give ALL the angles for which this is true, you would have to say pi/4 +k2pi and 3pi/4 +k2pi

OpenStudy (mertsj):

Do you understand?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

for my options i have a) pi/4 b) -pi/4 c) 0 d) -1 e) 1 i have to pick between 1 and -1...how do i decide

OpenStudy (mertsj):

Are we still talking about the same problem?

OpenStudy (anonymous):

yea

OpenStudy (mertsj):

Well then, I would go with 1 because I am assuming that you are a beginner in trig. If that right?

OpenStudy (mertsj):

Is that right?

OpenStudy (anonymous):

yea

OpenStudy (mertsj):

So go with 1

OpenStudy (anonymous):

ok kool. i have no way of telling whether its right or wrong. im just going over practice questions for a test but i understand the concept

OpenStudy (mertsj):

Good for you.

OpenStudy (anonymous):

Simplify sin 2k cos k + cos 2k sin k using an appropriate trigonometric identify. do you know this one?

OpenStudy (mertsj):

This looks like the formula: sin(x+y) = sin(x)cos(y)+ cos(x)sin(y) sin(2k)cos(k) + cos(2k)sin(k)

OpenStudy (mertsj):

It matches up so it would be sin(2k+k)=sin(3k)

OpenStudy (anonymous):

yea that looks right how about sinx + (cos^2x)/(sinx)

OpenStudy (mertsj):

Well, cos^2(x) = 1-sin^2(x) so if we make that substitution we would have sin(x) + 2-sin^2(x) /sin x

OpenStudy (mertsj):

Is the entire thing over sin(x) or just the cos^(x)?

OpenStudy (anonymous):

just the cos^(x) and we are simplyfing i forgot to mention that

OpenStudy (mertsj):